OpenStudy (anonymous):

What's the partial derivative w=rad(25-5x^2-5y^2)? The equation's longer, but I just need this piece of info cause I can't figure out the partial derivative of a radical

OpenStudy (turingtest):

use the chain rule just like in calc I, and treat the variables you are not differentiating with respect to as constants

OpenStudy (anonymous):

Does it stay in the radical, or does it become a fraction with the original equation in the denominator and the partial derivative in the numerator?

OpenStudy (turingtest):

let \(u=25-5x^2-5y^2\implies w=\sqrt u\), then \[\frac{\partial w}{\partial x}=\frac12u^{-1/2}\cdot\frac{\partial u}{\partial x}\]is that enough info?

OpenStudy (cruffo):

As @TuringTest said, let u = inside bit, then w' = u'/(2w)

OpenStudy (anonymous):

So..... \[\frac{ -10x }{ 2\sqrt{25-5x ^{2}-5y ^{2}} }\]?

OpenStudy (cruffo):

that's what I got, but you can simplify -10/2

OpenStudy (anonymous):

Ah, yes you can.... Alrighty, thanks!